# CoT-based Synthesizer: Enhancing LLM Performance through Answer Synthesis

### v1

@@ 5.4 Scaling Experiments
Figure 3: Effects of training data size on the performance of Scalar RM and Ours across the MATH and GSM8k datasets. As the data size increases, the performance of Scalar RM initially improves but subsequently declines while our method exhibits a consistent increase.
We evaluate the impact of training data size on both Scalar RM and our method. Specifically, we train the mathematical reasoning tasks with varying amounts of training data: 2k, 4k, 8k, 16k, 32k, 64k, 128k, keeping all settings consistent with the main experiment. Figure 3 illustrates the average performance trends of these two methods on the mathematical tasks as the data size increases.
@@
In contrast, Scalar RM exhibits a different trend: its performance initially with increasing data size but subsequently declines. This decline may be attributed to the limited training data diversity, which is generated through multiple samplings of the original benchmark dataset, leading to a significant number of duplicate instructions. Since Scalar RM is trained using positive and negative pairs and produces scalar scores during inference, it is more prone to “overfitting” on these repeated instructions. Additionally, as each dataset is trained for the same epoch, this overfitting becomes more pronounced with larger datasets. In comparison, our generative approach is inherently more robust to such repetition, remaining unaffected by the lack of diversity in the training data.
@@
We further investigate the impact of varying the number of candidate responses on the performance of policy models, including Self-consistency, ArmoRM, Scalar RM and our method. For this analysis, we use Llama3-8B-Instruct and Qwen2-7B-Instruct as policy models, conducting 125 inferences on the MATH test set. To ensure consistency across experiments, 1, 5, and 25 samples are randomly selected from the candidate answers for comparison.
Since our method involves concatenating multiple candidate responses before inputting them into the model, input length constraints prevent the model from processing all candidates simultaneously when the number of candidate responses becomes too large. To address this, candidate answers are grouped into sets of five, with any remaining answers treated as a separate group. The synthesized outputs from each group are then further combined iteratively across groups.
Figure 4: Effects of inference scaling on the performance of Scalar RM and Ours across the MATH and GSM8k datasets.
As shown in the Figure 4, as the number of candidate responses increases, the performance of our method and Self-consistency improves steadily, indicating their robustness. In contrast, the performance of the ArmoRM and Scalar RM, both of which adopt Best-of-N methods, shows an initial improvement followed by a decline as the number of candidate answers grows. This decline can be attributed to the verifier being distracted by other solutions containing erroneous information while searching for the optimal solution, leading to a “reward hacking” phenomenon. This observation aligns with the description provided in Cobbe et al. (2021).
@@ 5.5 Analysis
We select Llama3-8B-Instruct as the policy model and conduct a comparative analysis of SC, ArmoRM, Scalar RM and our method on the MATH test set. For each sample, the policy model generates 5 candidate responses, and we compare the number of correct answers as well as the final results after applying the post-processing methods, as shown in Table 3. Notably, our method can synthesize a correct answer even when all candidate responses are incorrect. Figure 5 provides a detailed example.
@@
Table 3: Distribution of correct answers for different post-processing methods on Llama3-8B-Instruct model across 5 inferences.
@@ 6 Conclusion
While our method supports the synthesis of multiple candidate answers, the input length limitation of the backbone model necessitates grouping the candidate answers for separate synthesis, followed by iterative synthesis of the grouped outputs. Although this approach leverages relationships among different candidate answers within the same group, it weakens the connections between candidates across different groups, and it also increases the overall inference cost. With the emergence of more advanced long-context LLMs, our method is expected to achieve more efficient and higher-quality solution generation by leveraging multiple candidate answers.
@@ Appendix A License
Appendix B Experiment Settings
B.1 Experimental Environment
B.2 Baselines and Implementation
The policy models are divided into two types, including API-based models and open-source models. For the open-source models that require us to deploy them ourselves, we use PyTorch, Transformers, and vLLM Kwon et al. (2023) to load the models. For each policy model, we generate five candidate responses. The inference parameters can be referred to in Table  4. For detailed information about each policy model, please refer to Table 6.
@@ B.2 Baselines and Implementation
For baseline methods, we first use the prompts specifically designed for each dataset to generate chain-of-thought (CoT) responses, with prompt details provided in Appendix C. The detailed information for the other baselines is introduced as follows:
B.3 Training Implementation
@@
Table 6: Information of baseline models.
@@
Appendix C Prompt for Inference
@@
Prompt for MATH CoT-prompt As an expert problem solver, solve step by step the following mathematical questions. Make sure that the last sentence is formatted as "The answer is XXX" and the value of the correct answer is surrounded by boxed when giving your answer. [Problem]: {question} [Solution]:
@@
Figure 5: An example of synthesizing a correct answer from all the wrong candidate responses generated by Llama3-8B-Instruct on MATH. The ellipsis indicates responses to the MATH instruction by CoT-prompting.

### v2

@@ 5.4 Scaling Experiments
Figure 3: Effects of training data size on the performance of Scalar RM and Ours across the MATH500 and GSM8k datasets. As the data size increases, Scalar RM initially improves but subsequently declines while our method exhibits a consistent increase.
We evaluate the impact of training data size on both Scalar RM and our method. Specifically, we train the mathematical reasoning tasks with varying amounts of training data: 2k, 4k, 8k, 16k, 32k, 64k, 128k. These subsets are sampled separately from the training datasets of Scalar RM and our method in the main experiment, while keeping all other settings consistent. Figure 3 illustrates the average performance trends of these two methods on the mathematical tasks as the data size increases.
@@
In contrast, Scalar RM exhibits a different trend: its performance initially improves with increasing data size but subsequently declines. This decline may be attributed to the limited diversity of training data, which is generated through multiple samplings of the original benchmark dataset, leading to a significant number of duplicate instructions. Since Scalar RM is trained using positive and negative pairs and produces scalar scores during inference, it is more prone to “overfitting” on these repeated instructions. Our generative approach, however, mitigates this issue by sampling multiple candidate answers for each question and synthesizing the final answer using CoT. Even when encountering repeated instructions, the sampling process ensures diverse candidate answers and reasoning chains, effectively transforming potentially repetitive training data into more diverse learning signals. This enables our approach more resilient to the lack of diversity in training data.
@@
We further investigate the impact of the number of candidate responses generated by the policy models on the performance of SC, ArmoRM, Scalar RM, and our method. Using Llama3-8B and Qwen2-7B as policy models, we conduct 125 inferences on the MATH500 test set, with 1, 5, and 25 samples randomly selected from the candidate answers for comparison.
Due to input length constraints, our method cannot process all candidate responses simultaneously when their number is large. To address this, candidates are grouped into sets of five, with any remaining answers treated as a separate group. The synthesized outputs from each group are then further combined iteratively across groups.
Figure 4: Effects of inference scaling on the performance of Scalar RM and Ours across the MATH500 and GSM8k datasets.
As shown in the Figure 4, the performance of our method and SC improves steadily with more candidate responses, indicating their robustness. In contrast, the performance of the ArmoRM and Scalar RM, both of which adopt Best-of-N methods, initially improve but then decline as candidate numbers grow. This decline can be attributed to the verifier being distracted by other solutions containing erroneous information while searching for the optimal solution, leading to a “reward hacking” phenomenon, which is consistent with the description in Cobbe et al. (2021). Additionally, regarding cost-benefit trade-offs, our method achieves comparable results with only 5 candidates, while SC requires over 10, demonstrating that SC needs significantly more candidates to match our method’s effectiveness. This underscores the efficiency and performance advantages of our method.
@@ 5.5 Analysis
We select Llama3-8B as the policy model and conduct a comparative analysis of SC, ArmoRM, Scalar RM and our method on the MATH500 test set. For each sample, the policy model generates 5 candidate responses, and we compare the number of correct answers as well as the final results after applying the post-processing methods, as shown in Table 3. Notably, our method can synthesize a correct answer even when all candidate responses are incorrect. Figure 5 provides a detailed example.
@@
Table 3: Correct answer distribution for different post-processing methods on Llama3-8B across 5 inferences.
@@ 6 Conclusion
While our method supports the synthesis of multiple candidate answers, the input length limitation of the backbone model necessitates grouping the candidate answers for separate synthesis, followed by iterative synthesis of the grouped outputs. Although this approach leverages relationships among different candidate answers within the same group, it weakens the connections between candidates across different groups, and it also increases the overall inference cost. Nevertheless, despite the additional inference overhead introduced by the input length constraint, the lightweight Synthesizer model remains more efficient and computationally economical than the policy model, which requires multiple inference steps. With the emergence of more advanced long-context LLMs, our method is expected to achieve more efficient and higher-quality solution generation by leveraging multiple candidate answers.
@@
Acknowledgments
This work is supported by the National Key Research & Develop Plan (2023YFF0725100) and the National Natural Science Foundation of China (62322214, U23A20299, U24B20144, 62172424, 62276270).
@@ Appendix A License
Appendix B Additional Ablation Experiments
Table 4: The impact of LLM Repair and Response LLM multiple sampling in ablation experiments of the data generation pipeline.
Policy Models
Method GLM4-api GPT-4o Llama3.1-70B Llama3.1-8B Llama3-8B Qwen2-7B Qwen2.5-14B Average
GSM8k
Synthesizer-8B(ours) 91.4 93.0 94.0 86.1 81.3 86.4 92.7 89.3
   w/o LLM Repair 90.6 -0.8 92.9 -0.1 93.6 -0.4 85.7 -0.4 81.5 +0.2 86.6 +0.2 92.0 -0.7 89.0 -0.3
   w/o multiple responses 90.5 -0.9 91.8 -1.2 93.4 -0.6 84.6 -1.5 80.8 -0.5 85.4 -1.0 91.8 -0.9 88.3 -1.0
MATH500
Synthesizer-8B(ours) 64.4 72.8 69.6 54.6 36.0 62.4 78.2 62.6
   w/o LLM Repair 63.6 -0.8 72.3 -0.5 68.4 -1.2 53.6 -1.0 35.2 -0.8 60.8 -1.6 77.8 -0.4 61.7 -0.9
   w/o multiple responses 63.4 -1.0 71.7 -1.1 68.0 -1.6 53.4 -1.2 33.6 -2.4 60.6 -1.8 77.0 -1.2 61.1 -1.5
Table 4 provides additional results on the impact of LLM Repair and Response LLM Sampling. The results show that both LLM Repair and Response LLM Sampling contribute positively to the overall performance of the Synthesizer-8B across different policy models and benchmarks. Removing LLM Repair leads to a slight drop in accuracy, while omitting Response LLM Sampling results in a more noticeable decline. This is because without Response LLM Sampling, the training data is limited to only 9k examples after filtering from the original 12k dataset. In contrast, removing LLM Repair still leaves 245k training examples, reducing the dataset by only 50k compared to the 295k used in the main experiment. This difference in dataset size explains the varying impact of the two ablation studies. Overall, incorporating both methods leads to more reliable and accurate responses.
Appendix C Experiment Settings
@@
Table 6: Information of baseline models.
@@
C.1 Experimental Environment
C.2 Baselines and Implementation
The policy models are divided into two types, including API-based models and open-source models. For the open-source models that require us to deploy them ourselves, we use PyTorch, Transformers, and vLLM Kwon et al. (2023) to load the models. For each policy model, we generate five candidate responses. The inference parameters can be referred to in Table  5. For detailed information about each policy model, please refer to Table 6.
Unlike traditional prompting methods that rely on iterative refinement, debate, or multi-step reasoning with decomposition and retrieval, our approach adopts an end-to-end synthesis strategy for efficiency and simplicity. Instead of introducing additional computational overhead, we directly generate a high-quality response in a single step by leveraging CoT reasoning to integrate information from multiple candidates. This streamlined design prioritizes efficiency, focusing on single-pass solution generation over multi-step reasoning.
To better evaluate the effectiveness of our approach, we further select the following baseline methods for comparison:
  • CoT-prompting: We use the prompts specifically designed for each dataset to generate CoT responses, with prompt details provided in Appendix E. CoT-prompting serves as a fundamental baseline, allowing us to assess the performance gains achieved by various post-processing methods.
C.3 Training Implementation
@@
Appendix D Model Definition
In this section, we provide detailed definitions and explanations of different LLMs for clarity. Specifically, in the data generation pipeline, the Sampling LLM and Response LLM are used to generate synthetic data, while the Policy Model and Base Model are employed during the Synthesizer’s inference stage, as illustrated in Figure 2. Below, we formally define these different LLMs:
  • Policy Model: In Synthesizer inference stage, it generates candidate responses to user queries. As shown in Table 1, the Policy Model can be either API-based (e.g., GPT-4o and GLM-4-Plus) or open-source LLMs (e.g., Llama3-8B-Instruct and Qwen2-7B-Instruct).
  • Base Model: In Synthesizer inference stage, it is used to train the Synthesizer. For simplicity, we use Llama3-8B-Instruct as the Base Model, which is also used as the Sampling LLM in our experiments.
  • Sampling LLM: In data generation pipeline, it generates diverse CoT candidate responses for queries. We use Llama3-8B-Instruct for this step. These candidates play a crucial role in subsequent processing.
  • Response LLM: In data generation pipeline, it synthesizes candidate responses into a final answer. We use Llama3.1-70B-Instruct for this step. The final correct answers will serve as training data. Our experiments show that even though the Synthesizer is trained with data from only a few models (Llama3-8B-Instruct, Llama3.1-70B-Instruct), it effectively enhances the performance of various Policy Models, indicating strong generalization capabilities.
Appendix E Prompt for Inference
@@
Prompt for MATH500 CoT-prompt As an expert problem solver, solve step by step the following mathematical questions. Make sure that the last sentence is formatted as "The answer is XXX" and the value of the correct answer is surrounded by boxed when giving your answer. [Problem]: {question} [Solution]:
@@
Figure 5: An example of synthesizing a correct answer from all the wrong candidate responses generated by Llama3-8B-Instruct on MATH500. The ellipsis indicates responses to the MATH instruction by CoT-prompting.